# Hadamard Representations: Augmenting Hyperbolic Tangents in RL

### v2

@@ 5 Related Work
Liu et al. (2019) investigated the need for sparse representations in the continuous control domain. Gulcehre et al. (2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues. Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022). Other work by Lyle et al. (2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays. Further research by Nikishin et al. (2022) used network resets to counteract the primacy bias and Sokar et al. (2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting. Nikishin et al. (2023) further studied plasticity injection for long-term training. In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024). This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
@@ 6 Conclusions and Discussion
This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem. Based on this analysis, we propose a novel representation architecture, the Hadamard Representation (HR), which enhances an encoder’s final hidden layer by taking the Hadamard product with a parallel, independently parameterized activation layer. We further discussed and empirically showed that applying the Hadamard representation with hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness. This approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension.
As investigated by Kumar et al. (2021), a compressed representation activated by a hyperbolic tangent attains a much higher effective rank than when activated by a ReLU, and will therefore yield stronger network expressiveness. In light of this, the hyperbolic tangent can potentially replace the ReLU in a variety of architectures where expressiveness is required without using deeper networks.
Future work could focus on further identifying the intricacies of hyperbolic tangents in an attempt to push the potential of continuously differentiable activations in reinforcement learning or even in a supervised learning setting (see Fig. 2). Also, we believe that an implementation of Hadamard-style architectures in a continual learning setting as in Abbas et al. (2023) could be promising.
@@ 7 Limitations
Our research has focused on the encoders’ final hidden layer in reinforcement learning. Both literature and preliminary experiments suggest that convolutional layers have a very strong preference for rectified activations (Jarrett et al., 2009), possibly due to a low-frequency bias (Ramasinghe & Lucey, 2022; Teney et al., 2024). We have therefore not conducted experiments integrating Hadamard representations in the convolutional section of our encoder architectures. In addition, a hyperparameter search could further strengthen our results, since the baseline is specifically tuned for the ReLU activation. Furthermore, as we perform experiments on DQN and PPO, integration of a Hadamard representation into more sophisticated algorithms and architectures such as Rainbow (Hessel et al., 2018) and Impala (Espeholt et al., 2018) would be interesting.
@@
Constructing a Hadamard representation is a relatively straightforward process, of which the Pytorch pseudocode can be found in Appendix A.1. The DQN and PPO algorithms used are adopted from the compact cleanrl baselines (Huang et al., 2022). Finally, the details on the KDE calculation and the effective rank calculation can be found in Appendix B.1 and Appendix B.2.
@@ C.2 Increasing Representation Parameters
Figure 14: Comparison of a normal hyperbolic tangent (Tanh), a hyperbolic tangent with a higher representation dimension $z_{t}\in\mathbb{R}^{512\rightarrow 1024}$ and a Hadamard representation using hyperbolic tangents. Comparisons are done on Performance (a), the fraction of dead neurons (b) and the effective rank of the representation $z_{t}$ (c). Naturally, increasing the representation dimension $z_{t}$ increases the effective rank of the representation, but using a larger representation dimension is not always preferable as it can negatively impact actual convergence.
@@ C.3 Validating dying neuron probability derivations
As discussed in Section 3, the effect of using a Hadamard representation strongly depends on the activation function. These derivations are empirically validated by the results in Fig. 15. In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1. This could explain the very slight increase in dead neurons when using an HR with activations.
@@ D.1 Evaluation Details
where Human Score and Random Score refer to the scores recorded by human players and random agents, respectively. Calculating our performance according to the Human-Normalized Score leads to the plot seen in Fig. 17. Due to taking a subset of the full atari domain, the VideoPinball environment is extremely dominant in the Human-Normalized Score calculation. For a more realistic comparison of the methods, we therefore decided to use baseline-normalized scores in the main paper.

### v3

@@ 5 Related Work
Liu et al. (2019) investigated the need for sparse representations in the continuous control domain. Gulcehre et al. (2022) analyzed network expressiveness in RL by measuring the effective rank (Kumar et al., 2021) of the compressed representation, and found that hyperbolic tangent representations generally maintain high rank while not suffering strongly from rank decay as training continues. Related work used normalization techniques and action penalization to counteract high variance in pixel-based robotic control (Bjorck et al., 2022). Other work by Lyle et al. (2022) investigated capacity loss in RL and similarly found that, as training progresses, the inherent network capacity of RL algorithms decays. Further research by Nikishin et al. (2022) used network resets to counteract the primacy bias and Sokar et al. (2023) evaluated and mitigated the dying ReLU phenomenon in DQN, both operating in the sample efficiency setting. Nikishin et al. (2023) further studied plasticity injection for long-term training and Delfosse et al. (2024) applied rational activations (Molina et al., 2019) in RL to increase plasticity. Concurrent work by Dohare et al. (2024) used continual back-propagation to further alleviate plasticity loss. In another related direction, recent work has investigated network sparsity in RL, showing that a large part of network capacity might be unnecessary when training reinforcement learning (Arnob et al., 2021; Graesser et al., 2022; Sokar et al., 2022; Tan et al., 2023; Obando-Ceron et al., 2024). This provides further insights into why a ReLU can achieve strong performance despite resulting in a significant number of dead neurons.
@@ 6 Conclusions and Discussion
This paper analyzed issues with continuously differentiable activations in RL and demonstrated that these activation functions also suffer from the dying neuron problem. Based on this analysis, we propose a novel representation architecture, the Hadamard Representation (HR), which enhances an encoder’s final hidden layer by taking the Hadamard product with a parallel, independently parameterized activation layer. We further discussed and empirically showed that applying the Hadamard representation with hyperbolic tangents reduces the occurrence of dead neurons in the representation and increases layer expressiveness. In DQN, PQN and PPO, this approach significantly improved performance in Atari games compared to both standard representation parameterizations and merely increasing the representation dimension. Finally, we have given insights into the ability of the ReLU activation to be less affected the symptoms of dying neurons, as opposed to non-zero saturating activations such as the hyperbolic tangent.
Future work could focus on further identifying the intricacies of the effects that different activation functions have on the resulting representation, in an attempt to push the potential of non piece-wise linear activations in reinforcement learning or even in a supervised learning setting (see Fig. 2). Also, we believe that an implementation of Hadamard-style architectures in a continual learning setting as in Abbas et al. (2023) or in Delfosse et al. (2024) could be promising.
@@ 7 Limitations
Using a Hadamard representation will double the incoming weights connected to a hidden layer. However, in Fig. 10, a comparison against the CReLU is shown, which equally doubles the network’s parameters. Furthermore, recent work shows that a scaling baselines in Atari often leads to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024). Second, due to the vast amount of activation functions and the theoretical sound candidacy for the hyperbolic tangent, we have mostly focused our research on the combination of hyperbolic tangents. Further research into multiplicative representations could find even better activation combinations. A hyperparameter search could also give more insights into the strengths of the Hadamard Representation, since the baselines are specifically tuned for the ReLU activation. Finally, as this research performs experiments on DQN, PPO and PQN, integration of a Hadamard representation into more complex algorithms and architectures such as Rainbow (Hessel et al., 2018), Impala (Espeholt et al., 2018) or SPR Schwarzer et al. (2021) would be interesting.
@@
Constructing a Hadamard representation is a relatively straightforward process, of which the Pytorch pseudocode can be found in Appendix A.1. The DQN and PPO algorithms used are adopted from the compact cleanrl baselines (Huang et al., 2022). The PQN algorithm uses the implementation in the Github repository of Gallici et al. (2024). Finally, the details on the KDE calculation and the effective rank calculation can be found in Appendix B.1 and Appendix B.2.
@@ A.1 Hyperparameters
Table 5: PQN Hyperparameters (Gallici et al., 2024)
Hyperparameter Value Description
Total Timesteps 10,000,000 Total timesteps for training
Timesteps for Decay 10,000,000 Timesteps for decay functions (epsilon and lr)
Number of Environments 128 Number of parallel environments
Steps per Environment 32 Steps per environment in each update
Number of Epochs 2 Number of epochs per update
Number of Minibatches 32 Number of minibatches per epoch
Starting Epsilon 1.0 Starting epsilon for exploration
Final Epsilon 0.001 Final epsilon for exploration
Epsilon Decay Ratio 0.1 Decay ratio for epsilon
Epsilon for Test Policy 0.0 Epsilon for greedy test policy
Learning Rate 0.00025 Learning rate
Learning Rate Linear Decay True Use linear decay for learning rate
Max Gradient Norm 10.0 Max gradient norm for clipping
Discount Factor $(\gamma)$ 0.99 Discount factor for reward
Lambda $(\lambda)$ 0.65 Lambda for generalized advantage estimation
Episodic Life True Terminate episode when life is lost
Reward Clipping True Clip rewards to range [-1, 1]
Frame Skip 4 Number of frames to skip
Max No-Ops on Reset 30 Max number of no-ops on reset
Test During Training True Run evaluation during training
Number of Test Envs 8 Number of environments used for testing
@@ C.2 Increasing Representation Parameters
(d) Learning Rates $z_{t}\in\mathbb{R}^{1024}$
Figure 15: Comparison of a normal hyperbolic tangent (Tanh), a hyperbolic tangent with a higher representation dimension $z_{t}\in\mathbb{R}^{512\rightarrow 1024}$ and a Hadamard representation using hyperbolic tangents. Comparisons are done on performance (a), the fraction of dead neurons (b), the effective rank of the representation $z_{t}$ (c) and learning rates of the higher-dimensional latent. Naturally, increasing the representation dimension $z_{t}$ increases the effective rank of the representation, but using a larger representation dimension is not always preferable as it often requires different hyperparameters, and can lead to reduced performance (Obando-Ceron et al., 2024; Obando Ceron et al., 2024).In (d), an ablation of learning rates shows that using a larger layer can sometimes prefer lower learning rates. However, it also shows that the improvement due to the Hadamard representation is likely not correlated with the parameter increase, as the Hadamard representation still significantly outperforms any of the 1024-dimensional latent state learning rate ablations.
@@ C.3 Validating dying neuron probability derivations
As discussed in Section 3, the effect of using a Hadamard representation strongly depends on the activation function. These derivations are empirically validated by the results in Fig. 16. In practice, since a neural network prefers symmetry, a sigmoid saturates slightly faster to 0 than to 1. This could explain the very slight increase in dead neurons when using an HR with activations. Note that, since we use neuron independence assumptions in our theoretically calculated dying neuron probabilities, the empirical results differ in magnitude from the theoretical predictions.
@@ C.4 ReLU activated Hadamard representation
C.5 Hyperbolic Tangent Convolutional Activation
We additionally run tests in DQN on 8 Atari games. The baseline ReLU is used, except for the convolutional activations which are changed to a hyperbolic tangent.
Figure 18: Baseline normalized performance in the Atari domain for 10M iterations (40M Frames) on DQN.
It is evident that using non-Hadamard hyperbolic tangents in the convolutional layers can lead to only a fraction of the performance of the ReLU-activated baseline.
@@ D.1 DQN Evaluation Details
where Human Score and Random Score refer to the scores recorded by human players and random agents, respectively. Calculating our performance according to the Human-Normalized Score leads to the plot seen in Fig. 19. Due to taking a subset of the Atari domain in DQN and PPO, the VideoPinball environment is extremely dominant in the Human-Normalized Score calculation. For a more realistic comparison of the methods, we therefore decided to use baseline-normalized scores in the main paper.
@@ D.2 Individual Environment Scores
Figure 22: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 23: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 24: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 25: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 26: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 27: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 28: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.